# Generalised Fermat equation: a survey of solved cases

### v1

@@ Acknowledgements
The authors sincerely thank Frits Beukers, Andrew Bremner, David Zureick-Brown, Tim Browning, Carmen Bruni, Pedro José Cazorla García, Andrzej Da̧browski, Rainer Dietmann, Jordan Ellenberg, Victor Flynn, Nuno Freitas, Andrew Granville, Alain Kraus, Florian Luca, Loïc Merel, Oana Padurariu, Kenneth Ribet, Jeremy Rouse, Samir Siksek, Sankar Sitaraman, Gökhan Soydan, Alain Togbé, Lucas Villagra Torcomian, Soroosh Yazdani, Konstantine Zelator and Zhongfeng Zhang for their responses, feedback and corrections on earlier versions of this paper. This research used the ALICE High Performance Computing facility at the University of Leicester.

### v2

@@ Acknowledgements
The authors sincerely thank Frits Beukers, Andrew Bremner, David Zureick-Brown, Tim Browning, Carmen Bruni, Pedro José Cazorla García, Andrzej Da̧browski, Rainer Dietmann, Jordan Ellenberg, Victor Flynn, Nuno Freitas, Andrew Granville, Alain Kraus, Florian Luca, Loïc Merel, Oana Pădurariu, Kenneth Ribet, Jeremy Rouse, Samir Siksek, Sankar Sitaraman, Gökhan Soydan, Alain Togbé, Lucas Villagra Torcomian, Soroosh Yazdani, Konstantine Zelator and Zhongfeng Zhang for their responses, feedback and corrections on earlier versions of this paper. This research used the ALICE High Performance Computing facility at the University of Leicester. We also thank the referees for their useful comments to improve the quality of this paper.